OpenStudy (anonymous):

5(x-6)-6(x-5)=x+5-(x-5)

OpenStudy (anonymous):

\[5(x-6)-6(x-5)=x+5-(x-5)\]

OpenStudy (anonymous):

I need help plz

OpenStudy (anonymous):

order of operations important here Brackets first! (5x-30)-6x+30 = x+5-x+5

OpenStudy (anonymous):

5x-6x-30+30=x-x+5+5

OpenStudy (anonymous):

-x = 10 x=-10

OpenStudy (anonymous):

thats your answer, x=-10